### 22 Dec 2023 v1

# Can Machines Learn Robustly, Privately, and Efficiently?

Abstract.
The success of machine learning (ML) applications relies on vast datasets and distributed architectures, which, as they grow, present challenges for ML. In real-world scenarios, where data often contains sensitive information, issues like data poisoning and hardware failures are common. Ensuring privacy and robustness is vital for the broad adoption of ML in public life. This paper examines the costs associated with achieving these objectives in distributed architectures.
We overview the meanings of privacy and robustness in distributed ML, and clarify how they can be achieved efficiently in isolation. However, we contend that the integration of these objectives entails a notable compromise in computational efficiency. We delve into this intricate balance, exploring the challenges and solutions for privacy, robustness, and computational efficiency in ML applications.
Byzantine robustness, federated learning, decentralized learning, differential privacy, cryptography
ccs: Computing methodologies Machine learningccs: Security and privacy Privacy-preserving protocolsccs: Security and privacy Distributed systems securityccs: Computing methodologies Distributed algorithms
1. Introduction
The landscape of computer science has been dramatically reshaped by the emergence of machine learning (ML). Initially conceived as a contender in simulated games (chess, jeopardy, go, poker, etc.), ML has evolved into an indispensable tool permeating diverse sectors such as finance and healthcare. In contemporary applications, ML algorithms power critical functions, from algorithmic trading in finance (Dixon et al., 2020) to the precise diagnosis of medical conditions (Sidey-Gibbons and Sidey-Gibbons, 2019). This evolution has been fueled by the confluence of abundant data and powerful computational architectures. Modern ML environments have shifted towards distributed settings for two compelling reasons. Firstly, the data, sourced from myriad locations, necessitates collaborative efforts where different entities combine their complementary datasets to train joint ML models. Secondly, the colossal computational tasks associated with the expansive ML models demand a distributed approach, deploying tasks such as gradient computations across multiple machines to expedite the training (Bertsekas and Tsitsiklis, 2015; McMahan et al., 2017).
However, the integration of distributed ML into sensitive applications presents critical trustworthiness challenges centered around two fundamental pillars: privacy and robustness. Failures in distributed frameworks, stemming from software bugs, hardware defects, or even deliberate perturbations, pose significant threats (Lamport et al., 1982) disrupting the algorithm in question. Ensuring that ML algorithms withstand corrupted training data and malicious machinations is therefore imperative in practical systems. Simultaneously, the sensitive nature of the data used in these applications necessitates stringent privacy safeguards. The exchange of model-related information among distributed components, essential for collaborative learning, can be exploited by adversaries, compromising the privacy of the underlying data (Zhu et al., 2019; Carlini et al., 2021).
Traditionally, the realms of robustness and privacy have been addressed in isolation in distributed ML, with separate strategies and techniques employed to enhance each of these two aspects. In practical systems, it is of paramount importance to address these fundamental issues in tandem. We explain in this paper that the confluence of privacy and robustness presents however a formidable challenge. While individual, efficient solutions have been proposed and empirically examined for both robust and private distributed ML (Kairouz et al., 2021; Allouah et al., 2023a), the direct amalgamation of these techniques to simultaneously ensure privacy and robustness is far from straightforward. As we shall explain, the merger of conventional privacy and robustness strategies in the context of distributed ML introduces a pivotal third dimension: efficiency. This dimension embodies the computational costs entailed by the proposed solutions. More specifically, we argue that safeguarding desirable performance or utility, often referring to accuracy in the context of machine learning, against strong and inquisitive adversaries necessitates a substantial escalation in computational resources. Nevertheless, we also provide insights into promising, computationally efficient alternatives that can deliver satisfactory accuracy for private and robust ML, when contending with weaker adversaries in terms of privacy and robustness.
A Case in Point: Income Census. To illustrate the complex interplay between robustness and privacy, consider the aggregation of annual incomes in a census. The objective here is to compute the average annual income of a population, under the dual constraints of ensuring (a) individual income privacy and (b) robustness against misreported data. In pursuit of privacy, individuals may introduce independent Gaussian noise to their true income figures, thereby obscuring these figures from curious entities, as in differential privacy (Dwork et al., 2014). Simultaneously, for the sake of robustness, the aggregation of reported incomes might be executed robustly, such as by calculating the median income (Yin et al., 2018), to mitigate the impact of grossly corrupted income data. Nonetheless, despite its computational efficiency, the introduction of noise to individual incomes inevitably leads to estimation errors (Steinke and Ullman, 2016; Allouah et al., 2023c). On the other hand, the robust aggregation, as exemplified by the median, diverges substantially from the true average when income distributions exhibit significant skew. A natural compromise for attaining both privacy and robustness is the robust aggregation of noisy incomes. Regrettably, this approach compounds inaccuracies, because the median does not attenuate individual noises, as averaging does, even within a large population. An appealing alternative surfaces in the form of homomorphic encryption, enabling the encryption of individual incomes instead of injecting noise. This approach permits the computation of the average income within the encrypted data domain, bypassing the need for preliminary decryption. While homomorphic averaging can be achieved efficiently (Zhang et al., 2020), it is not robust against corrupted income data. Furthermore, implementing non-linear robust aggregation methods, like the median, in a homomorphic manner presents substantial challenges and computational costs. This inefficiency becomes particularly pronounced when applied to high-dimensional inputs, a scenario commonly encountered in modern deep learning (Choffrut et al., 2023). This simple example serves as a microcosm illustrating the intricate relationship between privacy, robustness, and efficiency within the context of distributed ML.
This Paper. We first give a concise overview of the definitions of robustness and privacy in distributed ML, along with the techniques employed to attain these goals. We emphasize that privacy and robustness can be efficiently achieved independently, even against the strongest adversaries. We then shed light on the intricate three-way trade-off between privacy, robustness, and efficiency. Lastly, we explain that this trade-off can be mitigated for weaker threat models.
2. Threats and Guarantees
2.1. Training Architectures
Three main training setups have essentially been considered for distributed ML, as we depict in Figure 1. In centralized training, a trusted server collects data from individual machines, referred to as workers, performs the training, and subsequently distributes the resulting model back to these workers. This setting is rather simple, but it induces a high communication cost due to the sharing of the training data and relies on a trusted server, constituting a single point of failure. The federated setup employs a similar trusted server, but instead of sharing raw data, workers send gradients (or model parameters) computed on their datasets. These gradients are then averaged and utilized to iteratively update the model, akin to the approach used in Distributed Stochastic Gradient Descent (D-SGD) (Bertsekas and Tsitsiklis, 2015) and FedAvg (McMahan et al., 2017). This method effectively reduces communication costs. The decentralized setup is more resilient as it enables direct peer-to-peer communication among workers, alleviating the necessity for a central server. Each worker maintains a local model and aggregates information from peers through averaging gradients or parameters. Despite its resilience, decentralized training generally results in higher communication costs due to message redundancy among connected peers.
Figure 1. Three types of training setups for distributed ML. From left to right: centralized, federated, and decentralized.
2.2. Threats to Distributed ML
ML schemes need to face two crucial threats pertaining to robustness and privacy: poisoning and leakage, each posing distinct challenges, that also depend on the architecture of the distributed system.
Poisoning refers to the act of attacking a subset of workers, wherein the attacker corrupts the information transmitted by these workers to disrupt the training process. This corruption can target model-related information exchanged in both the federated and decentralized setups, such as gradients or models. Specifically, this form of attack manifests as model poisoning, orchestrated by Byzantine (a.k.a., active) workers (Blanchard et al., 2017). These adversaries are assumed to possess knowledge of the algorithm. They can collude and send meticulously crafted messages to the server or their peers, in the worst case. Additionally, poisoning can occur upstream in the ML algorithm, where the attacker corrupts the data of the workers (e.g., in centralized training), leading to what is known as data poisoning attacks (Biggio et al., 2012). While considered a somewhat weaker threat, it is very realistic and prevalent in practical scenarios (Shejwalkar et al., 2022).
Leakage refers to the ability of a curious adversary to deduce private information about the data held by the workers. This inquisitive (a.k.a., passive) adversary could be the workers, the server, or an external entity. For example, the curious adversary might attempt to discern the presence of specific data points in the training dataset by strategically querying the training model, as in membership inference attacks (Shokri et al., 2017). Other attacks may attempt to approximate the private data of the workers, as in data reconstruction attacks (Zhu et al., 2019; Carlini et al., 2021), where the adversary analyzes the gradients computed during training. For clarity, we exclude privacy threats originating from external entities, specifically those only capable of accessing the finalized published model. Our focus remains solely on threats arising from participants within the system during training time, encompassing other workers and, if present, the server. In other words, we consider the local model of privacy (Kasiviswanathan et al., 2011; Allouah et al., 2023c), which constitutes the strongest privacy threat model. Moreover, within the centralized and federated settings, it is common to assume the server to be honest-but-curious: it adheres to the protocol by correctly aggregating the received messages. Yet, the server may exploit the received information to compromise data privacy.
2.3. Trustworthy ML
To address these threats and establish trust in distributed ML, it is common practice to pursue Byzantine resilience for robustness and either differential privacy or cryptographic security for privacy. These are rigorous and widely studied standards of distributed ML, ensuring both robustness and privacy.
Robustness. Byzantine resilience aims to guarantee that the algorithm performs well under adverse conditions, even in the presence of faulty behavior from a fraction of the workers. In the context of the Byzantine ML literature (Blanchard et al., 2017; Yin et al., 2018; Farhadkhani et al., 2022; Allouah et al., 2023a), this translates to maintaining a low training error, even when a minority of workers exhibit Byzantine behavior. This assurance is typically conveyed through theoretical guarantees demonstrating the algorithm’s convergence to a favorable model, such as a stationary point or a local/global minimizer of the training loss (Farhadkhani et al., 2022; Allouah et al., 2023a). Notably, these guarantees do not rely on assumptions on the behavior or knowledge of Byzantine workers.
Privacy. Differential privacy (DP) (Dwork et al., 2014) aims to prevent a curious adversary from distinguishing between the outputs of the algorithm on two adjacent datasets, meaning datasets that differ by a single unit, e.g., data point in example-level DP, user in user-level DP (Dwork et al., 2014). This prevents a curious adversary from inferring the presence (or absence) of that specific data point in the dataset. DP is typically ensured by the injection of Gaussian or Laplacian noise in the algorithm, to the gradients or parameters. The degree of privacy in the resulting DP scheme is quantified by a positive number $\varepsilon$ which bounds the difference between the algorithm’s output distributions when given similar inputs, so the smaller $\varepsilon$ the more private the algorithm. In contrast, cryptography renders communicated messages unrecognizable to the adversary, typically via encryption. The level of privacy of the cryptographic security is measured by the length of the cryptographic key, which indicates the computational effort required by adversaries to breach the cryptosystem. The current security standards for homomorphic and symmetric encryption stand at 128 and 256 bits, respectively.
The choice of the training setup significantly impacts the achievable privacy and robustness goals. Centralized training is intrinsically not private, as workers’ data is sent directly to the curious server. [^1] Adding DP noise to data is an option, but it greatly decreases accuracy in general, except for very simple tasks (Fukuchi et al., 2017). The federated and decentralized setups are more suited to privacy, through the exchange of gradients or model parameters instead of raw data. Yet, recent attacks are able to reconstruct the training data from exchanged gradients (Zhu et al., 2019; Carlini et al., 2021), thus necessitating privacy-preserving solutions for these settings as well. In terms of robustness, centralized training also encounters challenges posed by Byzantine workers, but their influence is limited as these workers can only corrupt data. On the contrary, federated and decentralized training are more susceptible to Byzantine attacks. This vulnerability extends beyond mere data corruption to model poisoning attacks, impacting crucial aggregation operations essential to algorithms in these systems.
3. Privacy and Robustness Techniques
3.1. Privacy Techniques
Noise injection represents a pragmatic and cost-effective approach to ensure differential privacy (DP), as outlined in Dwork and Roth’s seminal work (Dwork et al., 2014). This technique essentially consists in introducing random noise, such as Laplacian or Gaussian noise, into the information communicated by the workers (e.g., gradients or models). In addition to being easily applicable across various topologies, this technique offers theoretical DP guarantees against curious adversaries equipped with any auxiliary information. However, for large-scale ML models, substantial noise may become necessary. Moreover, in iterative ML algorithms, the DP loss escalates with an increase in the number of iterations, yielding a well-studied tension between accuracy and privacy (Dwork et al., 2014; Steinke and Ullman, 2016).
Homomorphic encryption (HE) enables the execution of mathematical operations on encrypted data, eliminating the need for prior decryption. This technique yields results equivalent to conducting computations on plaintext data. By virtue of this approach, the confidentiality of workers’ messages is effectively protected, as it allows for the secure aggregation of encrypted gradients or parameters by the server and/or workers without direct visibility into their contents. Notably, HE finds its most prevalent application in the federated setting (Phong et al., 2018; Zhang et al., 2020; Ma et al., 2022). Furthermore, there is often a slight reduction in accuracy associated with HE solutions, primarily arising from the discretization of real values before encryption. This decrease is typically far less pronounced than the accuracy loss incurred due to DP (Zhang et al., 2020).
Secure multi-party computation (SMC) amalgamates encryption and oblivious transfer, offering a framework that enables multiple participants to collaboratively compute a shared function on their respective private inputs while preserving the confidentiality of these inputs. SMC is frequently employed in conjunction with HE primitives and has been extensively examined as a privacy-preserving solution in the context of distributed ML (Bonawitz et al., 2017; Fereidooni et al., 2021; Mohassel and Zhang, 2017). Although SMC has found common use in simpler ML models, such as linear and logistic regression, and decision trees (Liu et al., 2021), its adaptation to contemporary deep neural networks represents an ongoing challenge. Moreover, apart from the well-known bottleneck in communication complexity, it is worth noting that SMC-based implementations of ML algorithms involving more than two participants are known for incurring substantial computational overhead, especially when dealing with large datasets and number of participants (Mohassel and Zhang, 2017; Liu et al., 2021).
While other privacy-preserving techniques such as symmetric encryption and trusted execution environments (TEEs) have been considered, they exhibit some limitations. Symmetric encryption, though nearly costless today, is mostly relevant for the centralized and federated setups and relies on a strong assumption: the server is honest and non-curious. As for TEEs, they necessitate specialized hardware on both the server and workers, and they have demonstrated security vulnerabilities that sometimes compromise the privacy they are designed to protect (Fei et al., 2021).
3.2. Robustness Techniques
Robust aggregation emerges as a prevailing strategy when it comes to defending against Byzantine workers within federated and decentralized settings (Blanchard et al., 2017; Yin et al., 2018; Farhadkhani et al., 2022; Allouah et al., 2023a). It relies on non-linear aggregation techniques to counter adversarial influences, as any linear aggregation is arbitrarily manipulable by a single Byzantine worker (Blanchard et al., 2017). Namely, traditional averaging is substituted with robust aggregation methods, such as coordinate-wise median and trimmed mean (Yin et al., 2018), and Krum, Multi-Krum (Blanchard et al., 2017). This defense is executed at the server in federated learning or at the workers in decentralized learning, offering a robust barrier against Byzantine actors. Notably, these robust aggregation methods typically introduce minimal computational overhead, with some, like the coordinate-wise median, operating efficiently in nearly linear time in the number of aggregated inputs.
Suspicion-based defenses have been introduced, primarily in the federated setting, aimed at harnessing the assumption of a homogeneous data distribution across workers. These methods presuppose that the server possesses a proxy dataset, facilitating the detection of Byzantine workers. By assessing the alignment of worker gradients with the server’s reference gradient, or by employing martingale concentration arguments to monitor the temporal evolution of gradients, suspicion-based defenses seek to unmask adversarial actors (Alistarh et al., 2018; Xie et al., 2019). In controlled data center environments, it becomes possible to detect Byzantine workers using redundancy concepts borrowed from coding theory (Chen et al., 2018).
Data sanitization serves as a prevalent defense strategy in centralized training, when dealing with weaker adversaries only capable of corrupting a fraction of the data. This approach assumes that the server has direct access to the data, permitting the detection and removal of anomalous data points before subjecting the cleaned dataset to the learning algorithm. For instance, in classification tasks, data sanitization could involve setting a distance threshold to identify and exclude anomalous data points for each class (Steinhardt et al., 2017). An iterative data sanitization approach also exists, allowing the stepwise removal of anomalous data using feedback from the learning algorithm (Diakonikolas et al., 2019).
Among these defense strategies, robust aggregation stands out by offering robust protection against the strongest (i.e., Byzantine) adversaries with minimal assumptions. Robust aggregation methods have not only demonstrated strong convergence guarantees (Blanchard et al., 2017; Yin et al., 2018; Farhadkhani et al., 2022; Allouah et al., 2023a, b) but have also been validated empirically (Farhadkhani et al., 2022). While the other defense techniques may present cost-effective alternatives, they either assume a less potent adversary (as in data sanitization) or rely on assumptions that may not hold in practice, such as homogeneous data distribution or the presence of a trusted proxy dataset on the server. Lastly, it is essential to mention that when workers’ data exhibits extreme heterogeneity, there is an inherent robustness cost attributable to Byzantine actors, irrespective of the chosen defense mechanism (Karimireddy et al., 2022; Allouah et al., 2023b). Interestingly, the challenge of data heterogeneity can be tackled by adding pre-aggregation robust operations, e.g., NNM (Allouah et al., 2023a) and Bucketing (Karimireddy et al., 2022), aimed at reducing the effect of heterogeneity among honest workers.
{forest}
[…]
Figure 2. Techniques used to ensure privacy and robustness in distributed trustworthy ML.
4. Trade-off between Privacy, Robustness, and Efficiency
The challenge of upholding privacy and robustness in distributed ML inherently introduces computational efficiency as a crucial third dimension. When confronted with the stringent demands of DP or cryptographic security, in conjunction with Byzantine robustness, the pursuit of both privacy and robustness does not seamlessly integrate with computational efficiency. We argue that the academic community must invest more research effort in this domain to yield efficient solutions that can maintain satisfactory accuracy. In alignment with this objective, we highlight promising research directions that involve relaxing privacy concerns and/or mitigating the Byzantine adversary, potentially paving the way for a more computationally feasible convergence of privacy and robustness in distributed ML.
4.1. Efficiently Achieving Privacy or Robustness Separately
Privacy and robustness can be guaranteed separately with satisfactory computational overhead. Indeed, privacy and efficiency are simultaneously achievable by the means of noise injection. Ensuring DP via noise injection is computationally cheap, as it only requires the addition of random noise for each model parameter, which is linear in the size of the model $d$. Moreover, it induces a non-negligible privacy-accuracy trade-off. Specifically, for strongly convex distributed learning problems involving $n$ workers with $m$ data points per worker, this trade-off can be quantified as a training loss error in the order of $\tfrac{d}{\varepsilon^{2}nm^{2}}$, where $\varepsilon$ is the example-level DP budget (Steinke and Ullman, 2016; Allouah et al., 2023c). However, this trade-off is effectively amortized in systems with a large number of workers $n$ or data points $m$.
An alternative approach to efficiently ensuring privacy in distributed ML is HE, which upholds data privacy with cryptographic standards. For instance, Zhang et al. (2020) have introduced efficient modifications to the FedAvg algorithm (McMahan et al., 2017), enhancing its privacy defenses against curious adversaries within a federated setup by integrating the homomorphic Paillier cryptosystem. Zhang et al. (2020) report a low time requirement of $1.1$ seconds for homomorphic averaging (i.e., a single aggregation) at the server, even when dealing with a substantial model containing $1.25$ million parameters. This model was trained within a cross-silo federated setting involving nine workers. It is noteworthy that the solution proposed in (Zhang et al., 2020) only exhibits a marginal reduction in accuracy, specifically a mere 1% loss in accuracy. This reduction is attributable to the necessary quantization of real-valued gradients prior to encryption.
Similarly, when the consideration is solely on Byzantine adversaries and the privacy concerns are set aside, robustness and efficiency can be guaranteed simultaneously through the application of robust aggregation techniques. Our own experiments demonstrate the feasibility of executing a single aggregation of gradients using the robust median method, analyzed by Yin et al. (2018), in only $2\times 10^{-4}$ seconds. This evaluation was carried out on the same model and under the same experimental conditions as the study by Zhang et al. (2020). To provide a comparative baseline, we also conducted averaging operations on nine gradients of equivalent size, mirroring the conditions used in (Zhang et al., 2020). These trials yielded a mean execution time of $5\times 10^{-5}$ seconds. This outcome underscores the acceptable computational overhead associated with fortifying distributed ML algorithms using robust aggregation techniques.
4.2. The Computational Challenge of Privacy and Robustness
Differential Privacy and Robustness. A natural idea to efficiently ensure privacy and robustness is combining efficient techniques which guarantee these two objectives, when used in isolation. In fact, several works (Guerraoui et al., 2021; Zhu and Ling, 2022; Allouah et al., 2023c) have considered coupling noise injection for DP at the workers with robust aggregation methods at the server in the federated setting. However, Allouah et al. (2023c) have shown that there is a fundamental accuracy cost when simultaneously enforcing DP and Byzantine robustness. This accuracy cost predominantly arises from the inherent robustness cost associated with Byzantine workers in the presence of data heterogeneity (Karimireddy et al., 2022; Allouah et al., 2023b). In cases where the data across workers exhibits substantial heterogeneity, distinguishing the behavior of a minority of honest workers with outlier data from well-crafted malicious behavior becomes exceedingly challenging. Furthermore, broadly speaking, DP mechanisms introduce artificial data heterogeneity among workers through the injection of noise (Steinke and Ullman, 2016; Allouah et al., 2023c), thereby exacerbating the accuracy cost. Specifically, for strongly convex distributed learning problems involving $f$ Byzantine workers among $n$ workers, any Byzantine-robust and differentially private training algorithm must pay an additional training error scaling as $\tfrac{f}{n}\cdot\tfrac{1}{\varepsilon^{2}m^{2}}$ (Allouah et al., 2023c). In the worst case where there is a constant fraction of Byzantine workers in the system, this trade-off is not amortized even in systems with a large number of workers $n$, unlike the classical privacy-accuracy trade-off $\tfrac{d}{\varepsilon^{2}nm^{2}}$ (Allouah et al., 2023c). Yet, Allouah et al. (2023c) have introduced SMEA, a robust aggregation method that matches the aforementioned lower bound and attains the smallest training error among all known robust aggregation techniques. A noteworthy insight from SMEA is that it leverages the isotropic structure of the Gaussian noise used for DP to discard outliers. This contrasts with traditional robust aggregation methods, such as the coordinate-wise median and trimmed mean, which are not designed to work in conjunction with DP. However, SMEA is computationally inefficient, requiring an exponential runtime in the number of Byzantine workers (Allouah et al., 2023c). This emphasizes the complex trade-offs associated with combining DP and Byzantine-robust aggregation: one must choose between a traditional robust aggregation approach that offers reasonable execution times but comes with a considerable accuracy cost or SMEA, which theoretically provides tighter guarantees but incurs a considerable computational overhead.
Cryptographic Security and Robustness. In contrast, HE provides an effective means to sidestep the lower bounds associated with DP (Steinke and Ullman, 2016; Allouah et al., 2023c). HE schemes offer the ability to efficiently perform fundamental mathematical operations, such as averaging (Zhang et al., 2020). However, when dealing with more complex non-linear operations, e.g., median, nearest neighbors, or min/max operators, which are pivotal components of well-known Byzantine-robust aggregation techniques (Blanchard et al., 2017; Yin et al., 2018; Allouah et al., 2023a), the homomorphic implementation incurs considerable computational overhead (Choffrut et al., 2023). Hence, the seamless integration of HE with robust aggregation methods demands meticulous calibration before deployment. This critical consideration has been tackled by Choffrut et al. (2023), who investigated the most suitable robust aggregation method for homomorphic implementation in terms of computational efficiency. Choffrut et al. (2023)’s analysis advocates for the suitability of the coordinate-wise trimmed mean and median aggregations (Yin et al., 2018). These methods exhibit a remarkable alignment with batch-friendly HE cryptosystems, harnessing the advantages of batch processing. This facilitates the simultaneous execution of multiple clear-text operations as a single homomorphic computation, resulting in a substantial reduction in computational cost. This compatibility is significantly important, particularly when dealing with large deep learning models with millions of parameters. Despite the computational gains introduced by batching, the homomorphic implementation of the trimmed mean and median still grapples with efficiency challenges. Choffrut et al. (2023) report that the time required for a single robust aggregation in the homomorphic domain exceeds 2 minutes, even when applied to a model featuring 713,000 parameters in a federated topology involving nine workers. This represents a 110-fold slowdown in comparison to the 1.1-second execution time of its non-robust counterpart (Zhang et al., 2020), despite the consideration of a larger model in the latter work. Furthermore, it is crucial to emphasize that ML algorithms typically run for several hundred iterations, amplifying the significant computational workload associated with this approach. These quantitative findings, which underscore the “impracticality” of even the most promising Byzantine-robust aggregations in terms of homomorphic cost, emphasize the great challenges faced in designing ML algorithms that must concurrently defend against the strongest privacy and robustness adversaries.
4.3. Weaker Threat Models
So far, existing works have not been able to marry privacy, robustness, and efficiency against the strongest adversaries. Indeed, proposed solutions either suffer from a significant loss in accuracy via the use of DP (Zhu and Ling, 2022; Allouah et al., 2023c), or incur substantial computations when implementing robust aggregation methods using cryptographic primitives (Choffrut et al., 2023). Nevertheless, we argue that privacy and robustness can be efficiently combined when considering weaker adversaries. In general, in federated training, natural relaxations of the privacy threat model try to recover the honest non-curious server scenario, where the privacy-accuracy trade-off is known to be acceptable (Kairouz et al., 2021). Specifically, this trade-off is $n$ times smaller in theory, i.e., scaling as $\tfrac{d}{\varepsilon^{2}n^{2}m^{2}}$ for strongly convex learning problems, when trusting the server and protecting against a curious adversary only observing the global models (i.e., post aggregation at the server) during training. Furthermore, one can consider a federated-like setup with two non-colluding honest-but-curious servers to enable additive secret sharing techniques (Hao et al., 2021). These additional assumptions are also compatible with trust assumptions limiting the Byzantine attacks (as in suspicion-based defenses for instance), and enable more efficient solutions. For example, Hao et al. (2021) consider a federated-like setup where the two non-colluding servers possess proxy data for Byzantine detection. By doing so, they are able to train a model of $270,000$ parameters with $100$ workers in approximately $100$ seconds. The solution in (Hao et al., 2021) suffers a loss of at most $3\%$ in accuracy, when defending against a fraction of $20\%$ of Byzantine workers, compared to FedAvg (McMahan et al., 2017). One promising avenue for future exploration is the examination of the shuffle model (Cheu et al., 2019), which employs a trusted shuffler for worker identity anonymization, offering a relaxation of privacy threats with DP amplification guarantees. However, its study in the presence of Byzantine adversaries remains an unexplored research direction. Another interesting avenue for further research lies in weakening the conventional Byzantine threat model within the context of distributed ML. Instead of assuming that Byzantine workers consistently exhibit faults in every training iteration, one can consider an adversary who poisons workers’ models only with a certain probability (Yin et al., 2018). This weaker threat model aligns more closely with real-world practical systems and allows for the potential development of efficient HE-based solutions. These algorithms can capitalize on the existence of “optimistic” training iterations without faults, to sometimes perform homomorphic averaging instead of robust aggregation. This shift can significantly enhance the computational efficiency of the solution, at the cost of potentially weaker, probabilistic convergence guarantees.